A university reports that the middle 50% of the SAT 1 math scores of its students were between 575 and 670, with half the scores above 605 and har below. a. What SAT I math scores would be considered outliers for that university? Outliers would lie below or above b. What can you say about the shape of this distribution? The distribution is probably Show Work is REQUIRED for this question: Open Show Work
Answer #1

a. the definition of an outlier is any data point more than 1.5 IQRs below the first quartile or above the third quartile.

IQR = Q3-Q1 = 670-575 = 95

Outlier = Q1-1.5IQR & Q3+1.5IQR = [ 432.5,812.5]

b.

The distribution is Right skewed